### v1

# Is deeper always better? Replacing linear mappings with deep learning networks in the Discriminative Lexicon Model

Abstract
Recently, deep learning models have increasingly been used in cognitive modelling of language. This study asks whether deep learning can help us to better understand the learning problem that needs to be solved by speakers, above and beyond linear methods. We utilise the Discriminative Lexicon Model (DLM, Baayen et al.,, 2019), which models comprehension and production with mappings between numeric form and meaning vectors. While so far, these mappings have been linear (Linear Discriminative Learning, LDL), in the present study we replace them with deep dense neural networks (Deep Discriminative Learning, DDL). We find that DDL affords more accurate mappings for large and diverse datasets from English and Dutch, but not necessarily for Estonian and Taiwan Mandarin. DDL outperforms LDL in particular for words with pseudo-morphological structure such as slend+er. Applied to average reaction times, we find that DDL is outperformed by frequency-informed linear mappings (FIL). However, DDL trained in a frequency-informed way (‘frequency-informed’ deep learning , FIDDL) substantially outperforms FIL. Finally, while linear mappings can very effectively be updated from trial-to-trial to model incremental lexical learning (Heitmeier et al., 2023b, ), deep mappings cannot do so as effectively. At present, both linear and deep mappings are informative for understanding language.
Author’s note: This is a preprint. This paper has not been peer-reviewed.
1 Introduction
Over the last decade, deep learning models have been remarkably successful, culminating in the most recent large language models (e.g. OpenAI,, 2023; Anthropic,, 2023; Gemini Team,, 2023). Increasingly, deep learning has also made its way into cognitive modelling of language with studies exploring whether deep learning architectures can model aspects of language processing. Modelling efforts have been targeted at roughly four groups of models: predicting how humans form known and novel wordforms, such as inflections (e.g. rumelhart1986learning; Cardillo et al.,, 2018; Kirov and Cotterell,, 2018; Corkery et al.,, 2019; McCurdy et al.,, 2020), using activations in deep learning models to directly predict behavioural data or brain signals (e.g. Bhatia and Richie,, 2022; Salewski et al.,, 2023; Huang et al.,, 2024; Caucheteux and King,, 2022; Schrimpf et al.,, 2021; Hale et al.,, 2022; Goldstein et al.,, 2022), using deep learning components in theoretical models of lexical processing (e.g. Seidenberg and McClelland,, 1989; Harm and Seidenberg,, 2004; Magnuson et al.,, 2020; Scharenborg,, 2008), and finally, leveraging the iterative learning ability of deep learning models to model language acquisition (e.g. Plunkett et al.,, 1992; Li et al.,, 2007).
While these efforts have often been quite successful, using deep learning comes at the cost of ease of interpretation (but see e.g. Cammarata et al.,, 2020; Elhage et al.,, 2021; Bricken et al.,, 2023; Linzen and Baroni,, 2021). Deep learning models are universal function approximators, implying that if powerful enough, they can approximate any continous function with arbitrary accuracy (Hornik et al.,, 1989). For cognitive modelling, however, pure accuracy is not the end-goal: the question is whether increased precision afforded by deep learning models actually helps with predicting (cognitive) behaviour.
In the present paper, we address this question for the Discriminative Lexicon Model (DLM, Baayen et al.,, 2018, 2019), a model of (single) word comprehension and production. The model has been used to model the morphology of a wide range of languages (e.g. Chuang et al.,, 2020; Nieder et al.,, 2023; Heitmeier et al., 2023a, ; Chuang et al.,, 2023; van de Vijver and Uwambayinema,, 2022) but has also been successful at predicting behavioral data such as reaction times and acoustic durations (e.g. Schmitz et al.,, 2021; Stein and Plag,, 2021; Heitmeier et al., 2023b, ; Cassani et al.,, 2019; Chuang et al.,, 2021; Baayen and Smolka,, 2020; Saito,, 2023). Word forms are represented by high-dimensional binary vectors and semantics by high-dimensional real-valued vectors. In the DLM’s original formulation, comprehension is modelled with a linear mapping from word forms to word meanings and vice versa for production. Linear mappings are highly interpretable and surprisingly powerful, but also inherently limited by their linearity. In the present study we use the properties of linear and deep learning to investigated whether computational modeling can help us to better understand the learning problem that needs to be solved by speakers. Here, we are specifically interested in whether, and under what circumstances, deep learning generates insights above and beyond what more easily interpretable, linear models can tell us.
This paper is structured as follows: Section 2 briefly describes the DLM, the dense neural networks (termed Deep Discriminative Learning, DDL) we used to replace the linear mappings and how they were trained. Section 3 shows that DDL generally affords higher mapping accuracy, particularly on the training data, and explores for which words DDL outperforms LDL. Section 4 investigates DDL’s ability to predict lexical decision reaction times, both averaged across individuals and on a trial-by-trial basis. Finally, Section 5 discusses the results and concludes the paper.
2 Methods
2.1 Data
In the present study we used four datasets:
British Lexicon Project (BLP): Large-scale British English visual lexical decision database (Keuleers et al.,, 2012). The final dataset contained 28,482 word forms. We modelled this dataset with trigrams and word embeddings grounded in vision from Shahmohammadi et al., (2021).
Dutch Lexicon Project (DLP): Large-scale Dutch visual lexical decision database (Keuleers et al.,, 2010). The final dataset contained 13,675 word forms. We modelled this dataset with trigrams and fasttext embeddings (Grave et al.,, 2018).
Estonian: This dataset was first introduced in Chuang et al., (2020). The final dataset contained 4,093 wordforms of 230 lemmas, thus on average 17.8 wordforms per lemma (range from 2 to 28). We modelled this dataset with trigrams and fasttext embeddings.
Taiwan Mandarin: This dataset was first introduced in Chuang and Baayen, (2021) based on Fon, (2004). It contained 7,349 words with one to five syllables. We used a phone-based representation where vowels are combined with a number indicating lexical tone (e.g. {CJK*}UTF8gbsn 一下子, eng. abruptly, is represented as i1.x.ia4.z.ii5) and fasttext embeddings.
The four datasets pose different challenges: English and Dutch have limited inflectional morphology, therefore the challenge of the BLP and DLP datasets lies in modelling a large number of different lemmas. The dataset for Estonian, with its rich inflectional system, challenges our models to accurately map many different forms for only a few different lemmas. Finally, the challenge of the Mandarin dataset lies in modelling both segmental (phone) and suprasegemental (tone) information.
2.2 Models
In the Discriminative Lexicon Model (DLM), the simplest way to represent wordforms is by first splitting wordforms into letter/phone n-grams and then marking the presence and absence of n-grams in high-dimensional binary vectors. The form vectors of all wordforms in the lexicon are stored in a form matrix $\mathbf{C}$. Word meanings are likewise represented by high-dimensional vectors, usually taken from word embedding spaces such as word2vec or fasttext (Mikolov et al.,, 2013; Grave et al.,, 2018). They are then likewise stored in a semantic matrix $\mathbf{S}$. In Linear Discriminative Learning (LDL), comprehension is then modelled via a linear mapping $\mathbf{F}$ such that
$\hat{\mathbf{S}}=\mathbf{CF}$
and production via $\mathbf{G}$
$\hat{\mathbf{C}}=\mathbf{SG}.$
In Deep Discriminative Learning (DDL) we replace $\mathbf{F}$ and $\mathbf{G}$ with dense neural networks [^1] Note that we also experimented with LSTMs (Hochreiter and Schmidhuber,, 1997) and transformer architectures (Vaswani et al.,, 2017) but ultimately found that for the problem at hand simple dense neural networks provided the best generalisation accuracy while also being directly comparable to the linear setup.:
$\hat{\mathbf{S}}$ $=f(\mathbf{C})$
$\hat{\mathbf{C}}$ $=g(\mathbf{S})$
Please refer to Figure 1 for an overview of the number of hidden layers and dimensions for each model in this paper. We used Rectified Linear Units (ReLUs) as activation function throughout. The comprehension models were trained with a mean squared error loss. The production models additionally had a sigmoid activation function after the final layer and were trained with a binary cross-entropy loss. We used a batchsize of 512, except for the Estonian dataset where we used a batchsize of 256. Further training details depend on the experiment and are laid out in the sections below. More details on the training procedure can be found in Heitmeier, (2024).
(a) Comprehension
(a) Production
Figure 1: DDL model architectures. The red boxes represent dense layers (matrix multiplication plus bias) with the output dimension printed inside, blue boxes represent non-linearities. Box widths reflect the input and output dimensions of the layers. Figure from Heitmeier, (2024).
3 Does DDL improve mapping accuracy?
First, we need to verify that DDL does indeed improve accuracy of the mappings compared to LDL. Here, we focus both on performance on seen data as well as unseen data. We split the data into an approximate 80/10/10 split for training/validation/test data. The data was split in such a way that all letter n-grams in the validation and test data also occurred in the training data. We trained models for all four languages for a maximum of 500 epochs, with early stopping when validation accuracy did not increase for 20 epochs (except for the BLP production model, where we waited for 50 epochs since the model did not learn at all otherwise). As a baseline, we also trained LDL models on the joint training and validation data (since LDL does not require a validation set). We evaluated all models using correlation accuracy: if the predicted vector was more correlated with its target vector than with all vectors in the remaining data, it was counted as correct. We also computed accuracy@10 for the best models for each datasets. For accuracy@10, a mapping is counted as correct, if the target vector is among the 10 most correlated vectors. The results can be found in Figure 2.
(a) Comprehension
(b) Production
Figure 2: Correlation accuracy for DDL and LDL models for a British English, Dutch, Estonian and Taiwan Mandarin dataset. For DDL models, accuracies are averaged across 10 models, the error bar shows the standard deviation. DDL clearly outperforms LDL for Dutch and English, and less clearly for Estonian and Mandarin. Test@10 indicates the accuracy@10 on the test data for the best model.
It is clear that DDL outperforms LDL on both seen and unseen data for Dutch and English data where the challenge lies in mapping a diverse set of meanings. For Estonian and Mandarin on the other hand, LDL already performs surprisingly well and DDL only improves on LDL for production. However, the general increase in training accuracy across all four languages indicates that DDL does indeed afford higher mapping precision.
Next, we investigated for which words DDL outperforms LDL. We hypothesised that DDL would perform better than LDL for words which are harder to distinguish from other words. These are words with n-grams that are shared with many other words, shorter words and words in a dense semantic neighbourhood. We used these predictors (plus word frequency as a control) to predict the difference in correlation between predicted and target vectors in LDL and DDL using a Generalised Additive Model (GAM, Wood,, 2011; more details on how we operationalised these concepts can be found in Heitmeier,, 2024). Figure 3 shows that DDL indeed outperforms LDL for words with higher cue overlap (i.e. a word’s cues occur in many other words), shorter words and words in a denser orthographic neighbourhood (higher Coltheart’s N, Coltheart,, 1977). Moreover, it performs better on words with higher frequency. While semantic density (measured as the correlation between the target semantic vector and its closest semantic neighbour) seems to have a U-shaped effect, this is due to collinearity with the other predictors (Friedman and Wall,, 2005): in a GAM with only semantic density as a predictor, the measure shows the expected effect of an advantage for DDL for words in denser semantic neighbourhood.
Figure 3: Measures of between-word similarity predicting the difference between target correlation in DDL and LDL (target correlation DDL - target correlation LDL). For higher values of correlation difference, DDL outperforms LDL. We find that in general, DDL outperforms LDL for words that are similar to other words (i.e. words with higher cue overlap, a dense orthographic neighbourhood, shorter words and more frequent words). Figure from Heitmeier, (2024).
We also qualitatively examined words that showed particularly stark differences in correlation, favoring either LDL or DDL. The results can be found in Table 2. LDL outperforms DDL if words contain unique cues such as px or tv. DDL on the other hand shows a clear advantage for words which contain suffixes expressing a range of different meanings such as slender, barriers or choler which all contain trigrams which are often used in comparatives. Consequently, LDL places them in the semantic neighbourhood of other comparatives, while DDL maps them to their correct neighbourhood. This ability of DDL to work around the re-use of trigrams to express different meanings points to an interesting issue: language is full of re-use of sublexical features such as the -s ending expressing plurality (see also Shafaei-Bajestan et al.,, 2022; Plag et al.,, 2015), 3rd person singular, possessive or simply being a part of a word’s stem, or the -er ending which can express as varied meanings as comparative (e.g. hungrier), agent (e.g. teacher), instrument (e.g. poker), something associated with an action (e.g. diner) or denoting the origin of a person (e.g. Londoner) (Plag,, 2003, p.89). Any system modelling morphology needs to be able to deal with this “bricolage” (Lévi-Strauss,, 1962). Apparently, DDL is much more effective at doing so than LDL. This does not seem to be just an issue of precision: both Table 2 and the accuracies@10 in Figure 2 suggest that for LDL, the correct target word does not even make it to the top 10 most correlated words. A replication of this analysis for Dutch can be found in the Supplementary Materials.
Word Semantic neighbours LDL Semantic neighbours DDL Correlation difference
by by, standby, bye, written, thereby, brigades, dismissed, copied, outplayed, routed chastised, cajoled, bypassed, imposed, enforced, censured, challenged, embraced, shamed, entrenched -0.47
my my, her, me, your, own, myself, his, friend, our, dearest my, fathers, bruised, dearest, dad, childhood, dear, treasured, scarred, friend -0.41
px px, width, height, border, size, square, margin, padding, serif, background px, font, width, serif, size, apex, height, margin, cornered, tone -0.41
foursquare foursquare, twitter, mobile, hangouts, badges, twitted, yelp, badge, tweet, pollards foursquare, taproom, tavern, hangouts, beard, goatee, lapel, moustache, alehouse, stubble -0.35
tv tv, telly, cable, broadcasts, reruns, channels, airing, tuner, channel, viewers tv, telly, martin, shaw, john, cable, tom, mike, lightbulb, michael -0.34
(a) LDL outperforms DDL
Word Semantic neighbours LDL Semantic neighbours DDL Correlation difference
slang slashing, chopping, mangling, slamming, bashing, slating, slanting, hewing, chiding, slantwise slang, lingo, slangs, idiom, slur, patois, spelling, idioms, accent, rhymes 0.58
slender slighter, sturdier, trendier, rougher, livelier, finer, stouter, chunkier, bulkier, heartier slender, lithe, graceful, sinuous, lanky, tendrils, curving, threadlike, arched, leafless 0.58
charming arming, baiting, goading, harming, snaring, singeing, luring, robbing, maiming, gassing charming, quaint, demure, charm, classy, genteel, homely, urbane, serene, chatty 0.59
barriers livelier, chattier, sturdier, flashier, rowdier, trendier, sketchier, ruder, heartier, fleshier barriers, barrier, roadblocks, hurdle, curbs, impede, hindrance, crossings, hurdles, gaps 0.59
choler chunkier, wiser, moister, heartier, blunter, fattier, gunsmith, braver, plainer, trimmer choler, wherefore, foulness, rancour, straightway, withal, chequer, thine, vileness, smote 0.63
(b) DDL outperforms LDL
Table 2: Words for which LDL outperforms DDL the most (a) and for which DDL outperforms LDL the most. The two “semantic neighbours” columns show the words of the 10 semantic vectors which are most correlated with the vectors predicted by DDL and LDL respectively. The rightmost column shows the difference in target correlation between LDL and DDL. Table reproduced from Heitmeier, (2024).
4 Does DDL improve prediction of behavioural data?
Having ascertained that DDL does indeed generally improve mapping precision, particularly for the BLP and DLP, we were interested next in whether this also means that DDL outperforms LDL at predicting behavioural data. To this end we used visual lexical decision data from the BLP and DLP. In the following, we present results of predicting reaction times averaged across participants and predicting reaction times on a trial-by-trial basis following Heitmeier et al., 2023b .
4.1 Predicting average reaction times
For both the DLP and BLP we trained 5 comprehension models:
  • Endstate-of-Learning (EL): LDL model optimised equally across all words in the DLP/BLP.
  • Frequency-informed learning (FIL): LDL model trained in a frequency-informed way using a form of weighted regression (Heitmeier et al., 2023a, ).
  • Deep Discriminative Learning (DDL): DDL model trained on the 90% most frequent words in the DLP/BLP. Training was stopped after accuracy on the remaining 10% of words did not improve for 20 epochs.
  • Endstate-DDL (EDDL): DDL model trained on all words in the DLP/BLP for 2000 epochs to model a theoretical endstate of learning where all words are learned optimally.
  • Frequency-informed DDL (FIDDL): DDL model trained on the full token distribution of words in the DLP/BLP, i.e. the model was trained on each word as many times as its frequency count.
We used CELEX (Baayen et al.,, 1995) frequencies for both the BLP and DLP throughout. After training, we obtained the correlation between the predicted and the target semantic vectors for each model (“target correlation”). Reaction times were transformed as follows:
$\text{RTinv}=\frac{-1000}{\text{RT}}$
We used the target correlation measure to predict RTinv in one GAM per model. Target correlation was a highly significant predictor for reaction times across all models and both datasets ($p<0.0001$; summary tables for the GAMs can be found in the Supplementary Materials). Figure 4 shows the fit of the 5 models in terms of the Akaike Information Criterion (AIC, Akaike,, 1998) for both the DLP and BLP. First, we see that for both datasets, frequency-informed methods (FIL, FIDDL) outperform their non-frequency-informed counterparts (EL, DDL, EDDL). Regarding the difference between DDL and LDL, DDL and EDDL outperform EL for the DLP, but not for the BLP. Interestingly, while the correlation accuracy for EDDL was clearly higher than for DDL (DLP: 99% for EDDL compared to 92% for DDL; BLP: 100% compared to 88%), when it comes to predicting reaction times, EDDL only shows a slight advantage over DDL for the BLP and performs worse for the DLP. This indicates that better mapping accuracy does not necessarily entail better performance at predicting behavioural data. FIDDL outperforms all other models. From this we can conclude that while DDL does not necessarily outperform LDL, FIDDL clearly outperforms LDL both without and with frequency information. We will return to both these points in the final discussion.
(a) BLP
(b) DLP
Figure 4: Target correlation taken from an EL, FIL, DDL, EDDL and FIDDL model predicting reaction times in the BLP and DLP. Figure adapted from Heitmeier, (2024).
4.2 Predicting trial-to-trial reaction times
Next, we turn to using DDL models for predicting more fine-grained data, namely lexical decision data at the per-participant trial-to-trial level. Heitmeier et al., 2023b hypothesised that humans’ lexical knowledge is continuously changing and that this continuous change can be modelled with the DLM. They simulated the lexical decision experiment reported in the BLP. For each participant they ran one simulation where they updated the LDL mappings in the DLM using the Widrow-Hoff (Widrow and Hoff,, 1960) learning rule (dynamic simulation) and one where they did not update the mappings (static simulation). After applying one learning step of the Widrow-Hoff learning rule, the error between a predicted and target vector is slightly lower than before the update. This allows updating the mappings after each encountered stimulus in the lexical decision experiment in our simulations (details in Heitmeier et al., 2023b, ; Heitmeier,, 2024). Measures from these simulations were then used to predict individual reaction times in the BLP. They found that for a majority of subjects, dynamic simulations provided a better model fit to reaction times, particularly for words (as opposed to nonwords).
Here, we asked whether we can replicate these findings with DDL. To this end we repeated the analysis presented in Heitmeier et al., 2023b for the first 39 subjects in the BLP, replacing LDL mappings with DDL models and applying one backpropagation step (instead of the Widrow-Hoff learning rule) for updating the models after each trial. Details on the simulation procedure, the measures used for predicting reaction times and the statistical modelling can be found in Heitmeier, (2024).
We found that the AIC of all DDL-based models was worse than that of the LDL-based counterparts for all participants, for both words and nonwords and static/dynamic simulations (average AIC difference was -109.0/-131.6 AIC points for words for static/dynamic simulations and -126.1/-119.6 AIC points for nonwords for static/dynamic simulations). While for LDL we found that 72% of dynamic simulations showed a better fit than static simulations for words (33% for nonwords), for DDL 67% of dynamic simulations outperformed static ones for words (46% for nonwords). Figure 5 gives an overview over the average difference in AIC as well as the percentage of participants showing an advantage of trial-to-trial learning.
Figure 5: Comparison between static and dynamic simulations using DDL and LDL for predicting per-participant reaction times in the BLP. In general, the effect of dynamic simulations compared to static ones is larger for LDL, resulting in a bigger difference in AIC as well as a larger number of participants showing an advantage of trial-to-trial learning for LDL. Figure from Heitmeier, (2024).
When plotting the difference between the AICs of static and dynamic simulations with DDL against that with LDL (Figure 6), we found that they correlate for words, i.e. DDL and LDL generally agreed on which subjects show an effect of trial-to-trial learning as modelled by the dynamic simulation. No such correlation was present for nonwords.
(a) Words
(b) Nonwords
Figure 6: Comparison of AIC differences between dynamic and static simulations for DDL and LDL. Each dot represents one participant. The x-axis denotes the difference in AIC between dynamic and static simulations for the LDL models and the y-axis for the DDL models. Negative difference values indicate that the dynamic simulation outperforms the static simulation in terms of AIC, while positive difference values indicate the opposite. If a dot falls within a green shaded area, LDL and DDL agree on whether dynamic simulations provide a better model fit that static simulations for the respective participant. For words, the AIC differences are correlated ($r=.51,p=0.001$); for words no correlation is present ($p=0.436$). Figure from Heitmeier, (2024).
We conclude from this analysis that DDL is not able to model fine-grained trial-to-trial effects as well as LDL. DDL shows an advantage of learning for fewer subjects than LDL and the difference between dynamic and static is smaller for DDL than for LDL. However, at least for words, DDL and LDL largely agree on which subjects show an effect of trial-to-trial learning. A possible reason for the reduced effect of trial-to-trial learning in DDL is that a more accurate initial mapping may lead to less pronounced error feedback in the update step, thus each learning step having a smaller impact on subsequent trials. Moreover, the larger number of weights in the DDL networks might also mean that weight updates are distributed more evenly across weights and not centered as strongly on the weights of cues shared with subsequent words.
5 Discussion and Conclusion
In this paper we explored whether replacing the linear mappings (Linear Discriminative Learning, LDL) in the Discriminative Lexicon Model (DLM) with dense neural networks (termed Deep Discriminative Learning, DDL) improves the model’s mapping accuracy as well as its performance at predicting behavioural data. We found that across four languages (English, Dutch, Taiwan Mandarin, Estonian) DDL always improved training accuracy and also accuracy on held-out data for English and Dutch, the two datasets with the largest number of different lemmas. When predicting average lexical decision reaction times, we found that a frequency-informed variant of DDL (FIDDL) where the model is trained on each word according to its frequency outperformed all other methods (including frequency-informed LDL, FIL). However, when predicting fine-grained trial-to-trial lexical decision reaction times, DDL performs worse than LDL, even though the two generally agree as to which subjects show a better model fit when based on simulations where the DLM is dynamically updated after each trial.
It is worthwhile to focus on a few aspects of these results in more detail. First, we find a notable difference across datasets: In comprehension, DDL only reliably improves mapping accuracy over that of LDL for English and Dutch, but not for Mandarin and Estonian. This may to some extent be explainable by the respective data sizes: the English and Dutch datasets are larger (28,000 and 11,000 word forms) than the Estonian and Mandarin ones (6,000 and 3,000). However, the nature and diversity of the datasets may also play role. The English and Dutch datasets have a much larger number of different lemmas with few inflected word forms per lemma, while the Estonian dataset only has 230 lemmas with a large number of (relatively regular) inflected forms. For Mandarin, tone information contributes to a much larger number of cues (9,644, compare this to the much larger English dataset, which only has 7,030 cues), and as we saw in Section 3, fewer shared cues increase the performance of LDL models in comprehension. This suggests that the more semantically diverse a dataset, the more it profits from modelling with DDL. In production, on the other hand, all four models profit from DDL. In Heitmeier, (2024) we found that this can be accounted for to a significant extent by the loss function used: a linear model with the same number of trainable parameters as LDL but with added sigmoid + binary cross entropy resulted in almost the same accuracy as the deep learning models. However, the advantage of DDL over LDL in the present setup may also be due to the representations used. Chuang et al., (2024) found that LDL performs surprisingly well at mapping contextualised embeddings to pitch contours, and preliminary results suggest that in this setup, DDL does not improve over LDL even in production. This indicates that when both input and output are sufficiently discriminable, LDL suffices.
Second, notice that while DDL models outperformed LDL models in terms of pure mapping accuracy for both the English and Dutch datasets, they do not necessarily outperform LDL at predicting reaction times: DDL outperforms LDL at predicting average reaction times in the Dutch Lexicon Project (DLP), but not in the British Lexicon Project (BLP). The same is true for modelling lexical decision reaction times at the individual subject-level: even leaving aside trial-to-trial learning, the DDL models never perform better than the LDL models for static simulations. This highlights that pure mapping accuracy is not enough for predicting reaction times. However, it also leads to a further point, namely that modern machine learning and artificial intelligence may in some aspects be too good: human speakers make a lot of mistakes and are not particularly accurate. For example, participants in the DLP agree with themselves at chance level for low frequency words (below 10 per million) when tested on the same words twice (Diependaele et al.,, 2012). In the study of Ramscar et al., (2013), young adults (mean age 21.1 years) performed at chance level for low frequency words in an English lexical decision task (Balota et al.,, 1999). Moreover, the question arises of whether fully optimized performance in the DLM would be in any way helpful for predicting reaction times. Dense neural networks can in principle learn any continuous function to an arbitrary point of accuracy when given enough parameters (Hornik et al.,, 1989). In the case of the DLM this would mean that all predicted semantic and form vectors would be entirely accurate, thus losing any information we can gain from the DLM above and beyond simply using binary form vectors and pre-trained embeddings. The DLM’s imprecisions and mistakes offer valuable insights into possible processing bottlenecks.
Interestingly, from an evolutionary perspective, there is also some evidence that natural selection may not always favour the most optimal rule: Trimmer et al., (2012) found that in a simple evolutionary simulation, the Rescorla-Wagner learning rule was often favoured over other, more optimal learning rules, since it is more robust to variations in its parameterisation. For our setup this could suggest that even though deep learning models are more accurate, the higher number of parameters may make it less robust compared the simpler linear networks.
Third, examining the words that DDL maps more accurately than LDL is illuminating for linguistic analysis. We found that LDL outperforms DDL for words which contain very unique trigrams (e.g. #px, #tv), while DDL outperforms LDL when words contain trigrams which express a range of different functions such as -er. The re-use of building blocks is ubiquitous in language. While humans are able to deal with this “bricolage” relatively effortlessly, it is nevertheless a processing bottleneck which LDL struggles with. In this way, the comparison between LDL and DDL shows which components of single word processing may be particularly challenging. This finding could also have implications for theories of morphological priming (e.g. Taft and Forster,, 1975; Rastle and Davis,, 2008): our results suggest that LDL may place pseudomorphological words (e.g. slender, choler) in the semantic neighbourhood of morphologically complex words (e.g. slighter, chunkier). This could explain why morphologically transparent and morphologically opaque words are primed in a similar way (e.g. Creemers et al.,, 2020) even without morphological decomposition. Again, however, this result may, at least in part, be due to the representations used. An examination of the final segment of words ending in -er in the Buckeye corpus (Pitt et al.,, 2007) reveals that they have more varied pronunciations in nouns than in adjectives. These results dovetail well with the hypothesis that the nouns, which have much more varied meanings, have more differentiated realizations of the exponent — a distinction that is neutralized in the orthography. This in turn suggests that deep learning is required for -er words in English because the representations are impoverished, which opens the possibility that a linear mapping based on actual articulation might suffice.
Fourth, our results highlight the importance of the data distribution the models are trained on for cognitive modelling. Using a training/validation/test split is not a plausible mechanism for training cognitive models: after all, we usually encounter high frequency (seen) words, with the occasional low frequency (possibly unseen) word interspersed in-between. FIDDL (frequency-informed DDL) therefore has a much more plausible learning setup: it is trained on the full token distribution. By calculating the model’s accuracy across all words in the dataset we get a more realistic picture of how the model would fare with a real-world distribution of tokens (see also Heitmeier et al., 2023a, , for a discussion of token-based accuracy calculations). Surprisingly, this accuracy is very high (97% and 87% for Dutch and English respectively) even though the model saw 40% of the words in the English and 9% of the words in the Dutch dataset only 10 times or fewer. Such a setup also quells concerns about the performance of computational models on held-out data (Pinker and Ullman,, 2002). Humans are able generalise to unseen data, but in reality, word distributions are skewed towards high frequency words. Evidently, both humans and FIDDL perform excellently in such a setting. Moreover, these results once again highlight the importance of training models of lexical processing in a frequency-informed way when one is interested in a usage-based analysis of the data. The frequency-informed FIDDL and FIL models by far outperform their non-frequency informed counterparts when predicting average reaction times in the BLP and DLP.
To summarise, researchers interested in modelling a language system or behavioural data with the DLM may well profit from using DDL when they are working with a particularly large, diverse dataset which LDL struggles with and/or when they require a higher mapping accuracy for their modelling task. However, they should be aware that using DDL may not necessarily improve the model’s ability to predict behavioural data (unless trained in a frequency-informed way) and that DDL is costly: it requires experimenting with a number of hyperparameters as well as significantly more computational resources and training times.
Where does this leave us with regards to our initial question about deep learning’s role in understanding the learning problem which speakers face? Our comparison between LDL and DDL in the DLM lead us to a few conclusions. While deep learning models are generally more accurate, they are not a “tried-and-true” method for increasing cognitive modelling performance: pure accuracy improvements do not necessarily afford better performance at predicting behavioural data and may also not improve our understanding of the underlying processes. However, investigating cases where deep learning models clearly outperform linear models can nevertheless point to processing bottlenecks in a purely linear system which can be solved at the expense of additional parameters in deep learning models.
Acknowledgements
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC number 2064/1 – Project number 3907276455, and by the European Research Council, project WIDE-74254. The authors thank Yu-Ying Chuang, Tino Sering and Louis ten Bosch for helpful discussions.